# A Personalised Formal Verification Framework for Monitoring Activities of Daily Living of Older Adults Living Independently in Their Homes

### v1

@@ Abstract
There is an imperative need to provide quality-of-life to a growing population of older adults living independently. Personalised solutions that focus on the person and take into consideration their preferences and context are key. In this work, we introduce a framework for representing and reasoning about the Activities of Daily Living of older adults living independently at home. The framework integrates data from sensors and contextual information that aggregates semi-structured interviews, home layouts and sociological observations from the participants. We use these data to create formal models, personalised for each participant according to their preferences and context. We formulate requirements that are specific to each individual as properties encoded in Linear Temporal Logic and use a model checker to verify whether each property is satisfied by the model. When a property is violated, a counter example is generated giving the cause of the violation. We demonstrate the framework’s generalisability by applying it to different participants, highlighting its potential to enhance the safety and well-being of older adults ageing in place.
@@ 1 Introduction
Activities of Daily Living (ADLs) are essential self-care physical tasks that an individual performs on a regular basis. ADLs are fundamental for maintaining health and well-being [4]. Unexpected behaviour during these activities may indicate a decline in cognitive or physical abilities, potentially leading to unsafe conditions [5]. Moreover, monitoring such activities could be crucial for detecting changes that may impact an individual’s QoL. Sensors provide a powerful means of monitoring ADLs and, by extension, detecting deviations, thus potentially facilitating timely interventions that can safeguard or even improve the QoL of OAs.
In this paper, we present a novel framework that combines data from sensors, interviews, home layout and sociological observations (i.e. the systematic analysis of social behaviours and interactions) for representing and reasoning about ADLs in the homes of OAs living independently. The framework leverages unobtrusive, commercially-available sensors to capture ADL events, e.g. opening a refrigerator, closing a shower cubicle door. These events are integrated through an automation platform and stored in a time-series database. To enrich the sensor data, we systematically collect contextual information about daily routines through interviews and sociological observations. This contextual data also enables the customisation of sensor setups and the creation of accurate, person-centred models that account for home layout, privacy preferences, and individual requirements.
@@
By modelling the various actions involved in carrying out ADLs as a finite state model [6], we can form an abstraction of the behaviour of the OA. This enables the application of formal methods to verify whether the observed behaviour, captured by a set of sensors, satisfies a set of behavioural properties specific to that OA. These properties are derived from existing general guidelines and from specific requirements of the OA. We use model checking [6], a formal verification technique, to verify whether properties hold for the observed behaviour. More specifically, we use the model checker NuSMV [7] to specify and verify the model, encoding properties in a dialect of Linear Temporal Logic (LTL) [8].
@@
Our approach is generalisable and we have applied it to a diverse range of OAs. Here we illustrate with two participants with contrasting ADLs.
@@ 2 Related Work
Early work on ADL monitoring using unobtrusive sensors has primarily focused on recognising and tracking ADLs. Initial studies, such as those by Munguia et al. [11, 12], used contact sensors and energy switches to capture events involving people living alone to track daily activities. Following this, a range of studies have examined various aspects of ADL monitoring. For instance, Kang et al. [13] focused on the use of IoT devices and sensor fusion to capture ADL events and improve data reliability, while Matsui et al. [14] proposed semi-automated mechanisms that relied on physical buttons for activity annotation. Tonkin et al. [15] worked on a multi-sensor dataset collected in a controlled residential setting, enhancing its accessibility through improved documentation, annotations, and organisation. That dataset comprises short-duration scripted ADLs. The use of formal methods in ADL monitoring has also been explored in previous studies. Konios et al. [16] used Petri nets to distinguish between normal and abnormal behaviour in data captured from motion, light, temperature, and contact sensors. Garcia-Constantino et al. [17, 18] collected data from ambient and wearable sensors to identify abnormal behaviour during preparation of beverages in the kitchen using Petri nets. Magherini et al. [19] used temporal logic to automate the recognition of three ADLs within a hypothetical scenario involving a person living alone. The activities assumed specific sequences and timing of actions, which were captured by camera sensors.
A shared characteristic of the above approaches is their reliance on common assumptions about the ADLs a person performs and where these activities are conducted. Moreover, the aforementioned studies, involving formal verification and ADLs, were carried out in controlled, simulated environments. In contrast, our work focuses on OAs performing ADLs in their own homes and personalises the approach so that the data collection, models and properties to be verified are all tailored to their situation and preferences, an approach that has been advocated by other studies [20, 21]. Furthermore, we streamline our pipeline by using an automated tool that generates formal symbolic models based on home layouts and the positioning of sensors.
@@ 3.1 Framework Overview
Participants: Our study is a subject-centric approach. Through semi-structured interviews, participants provide insights into their daily routines, preferences and specific needs, which will later be expressed as LTL properties. This also allows us to personalise the placement of sensors, by considering individual needs, their living environment and other aspects that may otherwise be overlooked (see Section 3.2).
@@
Sensor Platform and Storage: We employ open-source solutions for sensor data acquisition and management. Home Assistant [23], a home automation platform, orchestrates data acquisition from our sensors. This provides a flexible and extensible setup for data aggregation. We store the sensor-generated data in InfluxDB [24], a high-performance time-series database optimised for handling large volumes of time-stamped data.
@@
Properties: For each participant we define a set of properties which will be formally checked. These properties are derived from existing safety guidelines and the individual’s contextual information, including needs and preferences , creating requirements for the expected behaviour. We encode the properties in LTL and use a model checker to verify that the model satisfies the specified requirements (see Section 4.3).
@@ 3.2 Data Collection Process
In what follows, we describe the qualitative and sensor data collection. For different participants, data are collected over periods of one to three weeks depending on their availability.
@@ 3.2.1 Qualitative Data
Once consent is obtained, a home visit is scheduled with the participant for the collection of qualitative data. A semi-structured interview is used by the research team to gauge the participant’s daily routines, the context in which these are performed and the needs and preferences the participant may have prior to sensors installation. The interview guide is divided into five sections: i) ADLs, ii) context and environment including the impact of sensors on the daily routine, iii) thoughts on privacy and security, iv) current needs and preferences, feedback and suggestions and v) general well-being. From the interview and observations, we also obtain insight into the possible placement of sensors. For example, some participants express discomfort with motion sensors in private areas such as bedrooms, but are comfortable with contact sensors in the same room. Figure 2a shows an example of the layout and the placement of sensors for one participant, based on their activities and privacy concerns. At the end of sensor data collection, a second interview is conducted to assess any changes in the participant’s routines or needs and to determine whether their behaviour was influenced by the presence of sensors. Qualitative data are used to identify requirements that can be translated into properties regarding the participant’s behaviour.
@@ 3.2.2 Sensor Data
Sensor data are collected from contact and motion sensors placed throughout the participant’s home. The latter are positioned to cover living spaces and areas of transit, while contact sensors are attached to objects like cupboard doors or drawers. Over the course of our study, the number of sensors installed in homes ranged from to , while data collection ranged from to days. Passive infrared (PIR) sensors are used detect movement, with a detection angle of and a range of up to metres. These sensors have two possible states: clear and detected. If motion is registered while the sensor is in a clear state, the state changes to detected and remains in that state for seconds. If no further motion is registered within seconds, the sensor switches back to the clear state. If motion is registered while the sensor is in the detected state, the 30-second timer is restarted. Each contact sensor has two possible states: open and closed. The contact sensor consists of two components: body and magnet. When the body is within mm of the magnet, the state of the sensor is closed and beyond this distance, the state is open. Additionally, each contact sensor registers its internal temperature. Figure 3a shows a motion sensor used to detect movement in the area where medication is kept on a shelf in a participant’s kitchen. Figure 3b shows contact sensors attached to various drawers in the participant’s kitchen. Both motion and contact sensors operate on an event-driven basis and are connected to a ConBee II Zigbee gateway. As previously mentioned, we use Home Assistant to orchestrate data acquisition from the sensors and store our data using InfluxDB. The automation platform, gateway and database run on an NVIDIA Jetson Nano with a quad-core ARM CPU, a -core GPU, and GB of RAM, operating on a Linux-based system. The collected data are exported from the database in CSV format, resulting in dataset sizes ranging from MB to MB.
@@ 4.1.1 Information from participants
The data generated from the interviews allow us to learn about each individual’s ADLs, personal preferences and living context, which, in turn, help us model the properties that reflect general and specific ADLs. We apply Framework Analysis [25], a qualitative analysis method, to reveal patterns and themes in the interview data using NVivo software (Limivero, v14) [26]. We identify properties of ADLs by considering both our questions during interviews and the routines participants highlighted as important. These properties are then categorised as general (common to everyone, like eating or washing), individual-specific (like pet care or hobbies), or routine but non-daily activities that impact health (like using medication). This approach allows us to have a detailed representation of the properties used to verify general and specific ADLs.
@@ 4.1.2 Sensors
A contact sensor broadcasts its current state every time it changes from open to closed, closed to open or when there is a fluctuation of in the device’s temperature. Repetitions of the same state are taken to be noise in the data triggered by temperature fluctuations so we remove consecutive open (resp. closed) states and consider only the first occurrence. This provides a more accurate representation of the participant’s interaction.
A motion sensor broadcasts its state whenever there is a transition between clear and detected states. If the state is detected and no additional motion is registered within seconds, the sensor broadcasts an artificial clear state after the seconds. We remove clear states from the data as they do not reflect the actual cessation of motion. This provides a more accurate representation of movement, treating motion sensor events as instantaneous. In addition, if continuous clear and detected states occur for a motion sensor, with no changes in other sensors, we only consider the initial detected state of the motion sensor and discard the subsequent states.
@@ 4.2 Model generator
We use NuSMV to specify a model that represents the home and behaviour of the participant. The model is a transition system where changes in sensors can occur one at a time. This is in line with what we observe in the data. The model consists of: i) states, which represent possible configurations of sensor states and where the participant is during the execution of the system; ii) transitions, that correspond to events, which describe how the system evolves from one state to another; and iii) the initial conditions that define the sensors’ starting configuration. The model is verified against a set of logically-encoded properties (see Section 4.3).
@@
| Motion | Corridor | Facing corridor | MotionCorridor |
@@
The behaviour of each sensor in the model is encapsulated in a module. An excerpt of the NuSMV code for the contact module is shown in Listing 1. For each contact sensor we define three variables: its state, activation and deactivation, which are initialised as FALSE (see lines - ). In the module, the state of a contact sensor is a boolean variable where FALSE indicates the body and magnet of the sensor are nearby (e.g. drawers closed as shown in Figure 3b) and TRUE indicates they are apart. The activated boolean variable is set to TRUE at the time step when the state transitions from FALSE ( state) to TRUE (state), as shown in line , and it remains TRUE for one time step only. Note that while activated represents interaction with the person and thus serves as evidence of their activities, the state of a sensor being open – which can last for more than one time step – does not. The deactivated boolean variable is set to TRUE at the time step when the state transitions in the opposite direction, as shown in line , also remaining TRUE for one time step only. Similarly, while deactivated represents interaction by a participant, the state of a sensor being closed does not.
@@
We provide the model with a subtrace of events that were actually captured by the different sensors. This subtrace represents the observed behaviour of a participant during a specific period of time. Consider a trace for a set of sensors . Given a pair of unique time markers, and – manually inserted to mark the start and end of a time interval – the subtrace is the set of all events, chronologically ordered, from to , where denotes the time of an event:

### v2

@@ Abstract
There is an urgent need to provide quality-of-life to a growing population of older adults living independently. Solutions that focus on the person and take into account their preferences and context are recognised as key. We introduce a framework for representing and reasoning about the Activities of Daily Living of older adults living independently at home. The framework integrates data from sensors and data from participants derived from semi-structured interviews, home layouts and additional contextual information, such as the researchers’ observations. These data are used to create formal models, personalised for each participant according to their preferences and context. Requirements specific to each individual are formulated and encoded in Linear Temporal Logic, and a model checker is used to verify whether each is satisfied by the model of the participant’s behaviour. We demonstrate the framework’s generalisability by applying it to two different participants, highlighting its potential to enhance the safety and well-being of older adults ageing in place.
@@ 1 Introduction
Activities of Daily Living (ADLs) are essential self-care physical tasks that an individual performs on a regular basis. ADLs are fundamental for maintaining health and well-being [4]. Unexpected behaviour during these activities may indicate a decline in cognitive or physical abilities, potentially leading to unsafe situations [5]. Moreover, monitoring such activities could be crucial for detecting changes that may impact an individual’s QoL. Sensors provide a powerful means of monitoring ADLs and, by extension, detecting deviations, thus potentially facilitating timely interventions that can safeguard or even improve the QoL of OAs.
In this paper, a novel framework is presented that combines data from sensors, interviews, home layout and sociological observations (i.e. the systematic analysis of social behaviours and interactions) for representing and reasoning about ADLs in the homes of OAs living independently. The framework leverages unobtrusive, commercially-available sensors to capture ADL events (e.g. opening a refrigerator, closing a shower cubicle door). These events are integrated through an automation platform and stored in a time-series database. To enrich the sensor data, contextual information about daily routines is systematically collected through interviews and sociological observations. This contextual data also enables the customisation of sensor setups and the creation of accurate, person-centred models that account for home layout, privacy preferences, and individual requirements.
@@
By modelling the various actions involved in carrying out ADLs as a finite state model [6], we can form an abstraction of the behaviour of the OA. This enables the application of formal methods to verify whether the observed behaviour, captured by a set of sensors, satisfies a set of behavioural properties specific to that OA. These properties are derived from existing general guidelines and from specific requirements of the OA. Model checking [6], a formal methods technique, is used to verify whether the properties hold for the observed behaviour. More specifically, the model checker NuSMV [7] is used to specify and verify the model, encoding properties in a dialect of Linear Temporal Logic (LTL) [8].
@@
Our approach is generalisable and has been applied to a diverse range of OAs. Here we illustrate it with two participants with contrasting ADLs.
@@ 2 Related Work
Early work on ADL monitoring using unobtrusive sensors has primarily focused on recognising and tracking ADLs. Initial studies, such as those by Munguia et al. [11, 12], used contact sensors and energy switches to capture events involving people living alone with the aim of tracking daily activities. This was expanded upon by teams deploying sensors in the homes of OAs to monitor ADLs specifically for health care purposes. Rantz et al. [13] used temperature, motion, pulse and respiration sensors to generate alerts notifying health care providers of potential illness or functional decline. Kaye et al. [14, 15] used sensor systems in the homes of OAs to capture health related metrics as well as other measures of interest, such as mobility and sleep behaviour, in real-time. Their subsequent work [16] outlined a protocol for sensor deployment in homes focused primarily on technical aspects of installation and system functionality. While such technical rigour is essential, optimal sensor placement also requires user-centred approaches that consider individual routines and preferences. Skubic et al. [17] developed a continuous in-home monitoring system for OAs using unobtrusive motion and bed sensors. Their system generates alerts when statistically significant changes indicate a potential illness or functional decline. More recently, Janes et al. [18] discussed the use of in-home sensor systems combined with health records to slow functional decline in people living with lateral sclerosis. Similarly, Sprint et al. [19] explored how unobtrusive home sensors can be leveraged to monitor daily activities and assess cognitive impairments.
While the aforementioned work, frequently conducted with clinical support, has primarily focused on health care, our work complements this perspective by addressing the quality of life of OAs by verifying that their routines are performed in accordance with safety and behavioural properties in combination with their own preferences. Building on this distinction, subsequent studies have explored various aspects of ADL monitoring beyond health care contexts. For instance, Kang et al. [20] focused on the use of IoT devices and sensor fusion to capture ADL events and improve data reliability, while Matsui et al. [21] proposed semi-automated mechanisms that relied on physical buttons for activity annotation. Tonkin et al. [22] worked on a multi-sensor dataset collected in a controlled residential setting, enhancing its accessibility through improved documentation, annotations, and organisation. That dataset comprises short-duration scripted ADLs.
@@
The use of formal methods in ADL monitoring has also been explored in previous studies. Konios et al. [23] used Petri nets to distinguish between normal and abnormal behaviour in data captured from motion, light, temperature, and contact sensors. Garcia-Constantino et al. [24, 25] monitored ambient and wearable sensor data to identify deviations from normal behaviour during preparation of kitchen beverages, also using Petri nets. Magherini et al. [26] used temporal logic to automate the recognition of three ADLs within a hypothetical scenario involving a person living alone. The activities assumed specific sequences and timing of actions, which were captured by camera sensors. A shared characteristic of the above approaches is their reliance on common assumptions about the ADLs a person performs and where these activities are conducted. Moreover, these studies, involving formal verification and ADLs, were carried out in controlled, simulated environments.
As discussed by Kessler et al. [27] and Alexander et al. [28], a person-centred approach is increasingly recognised as important when recommending and using technology to monitor OAs’ daily activities. Other examples of person-centred tailoring in different domains include work on activity recognition by Guerra et al. [29] and on supporting better self-management of health information for OAs by Robinson et al. [30]. In our work, we focus on OAs performing ADLs in their own homes and personalise the approach so that the data collection, models and properties to be verified are all tailored to their situation and preferences.
@@ 3.1 Framework Overview
Participants: Our study is a person-centric approach. Through semi-structured interviews, participants provide insights into their daily routines, preferences and specific needs, which will later be expressed as LTL properties. This also allows us to personalise the placement of sensors, by considering individual needs, their living environment and other aspects that may otherwise be overlooked (see Section 3.2).
@@
Sensor Platform and Storage: We employ open-source solutions for sensor data acquisition and management. Home Assistant [32], a home automation platform, orchestrates data acquisition from our sensors. This provides a flexible and extensible setup for data aggregation. The sensor-generated data are stored in InfluxDB [33], a high-performance time-series database optimised for handling large volumes of time-stamped data.
@@
Properties: For each participant we define a set of properties which will be formally checked. These properties are derived from existing safety guidelines and the individual’s contextual information, including needs and preferences, creating requirements for the expected behaviour. The properties are encoded in LTL and a model checker is used to verify that the model satisfies the specified requirements (see Section 4.3).
@@
To streamline the pipeline, an automated tool is used to generate formal symbolic models based on home layouts and the positioning of sensors.
@@ 3.2 Data Collection Process
In what follows, we describe the qualitative and sensor data collection. For different participants, data are collected over periods of one to three weeks. This duration was the result of a number of factors within the six-month period allocated for sensor deployment. Installations were timed to avoid visitors, minimising noise caused by the presence of another person and ensuring data reflected typical ADLs. In addition, participants’ plans (e.g. hospitalisation, relocation, holidays) further constrained the deployment period. Another factor was the availability of the team of researchers carrying out and managing the concurrent interviews and sensor deployments.
@@ 3.2.1 Qualitative Data
Once consent is obtained, a home visit is scheduled with the participant for the collection of qualitative data. A semi-structured interview is used by the research team to gauge the participant’s daily routines, the context in which these are performed and the needs and preferences the participant may have prior to sensors installation. The interview guide is divided into five sections: i) ADLs, ii) context and environment including the impact of sensors on the daily routine, iii) thoughts on privacy and security, iv) current needs and preferences, feedback and suggestions and v) general well-being. From the interview and observations, we also obtain insights into the possible placement of sensors. For example, some participants express discomfort with motion sensors in private areas such as bedrooms, but are comfortable with contact sensors in the same room. Figure 2a shows an example of the layout and the placement of sensors for one participant, based on their activities and privacy concerns. At the end of sensor data collection, a second interview is conducted to assess any changes in the participant’s routines or needs and to determine whether their behaviour was influenced by the presence of sensors. Qualitative data are used to identify requirements that can be translated into properties regarding the participant’s behaviour.
@@ 3.2.2 Sensor Data
Sensor data are collected from contact and motion sensors placed throughout the participant’s home. The latter are positioned to cover living spaces and areas of transit, while contact sensors are attached to objects like cupboard doors or drawers. Passive infrared (PIR) sensors are used to detect movement, with a detection angle of and a range of up to metres. These sensors have two possible states: clear and detected. If motion is registered while the sensor is in a clear state, the state changes to detected and remains in that state for seconds. If no further motion is registered within seconds, the sensor switches back to the clear state. If motion is registered while the sensor is in the detected state, the 30-second timer is restarted. Each contact sensor has two possible states: open and closed. The contact sensor consists of two components: body and magnet. When the body is within mm of the magnet, the state of the sensor is closed and beyond this distance, the state is open. Additionally, each contact sensor registers its internal temperature. Figure 3a shows a motion sensor used to detect movement in the area where medication is kept on a shelf in a participant’s kitchen. Figure 3b shows contact sensors attached to various drawers in the participant’s kitchen. Both motion and contact sensors operate on an event-driven basis and are connected to a ConBee II Zigbee gateway. As previously mentioned, Home Assistant is used to orchestrate data acquisition from the sensors and the data are stored using InfluxDB. The automation platform, gateway and database run on an NVIDIA Jetson Nano with a quad-core ARM CPU, a -core GPU, and GB of RAM, operating on a Linux-based system. The collected data are exported from the database in CSV format, resulting in dataset sizes ranging from MB to MB.
Data were collected from sensors deployed in homes, with to sensors per home and data collection periods of to days. There was no restriction on the number of sensors deployed in each home. Sensor placement and type were guided by qualitative analysis of participants’ routines and home layouts. Participants were older adults ( women, men) aged – years, recruited through community centres and library postings. The participants lived in areas ranging from the th to the th decile of the Scottish Index of Multiple Deprivation (SIMD) [34]. SIMD is a standard measure of relative deprivation across small geographic areas in Scotland, combining indicators such as income, health and crime. This indicates that all participants resided in areas from mid to least deprived. For this work, we present data from two participants from the 5th and 7th deciles, one male and one female, to demonstrate the approach rather than report participant-specific results. The evaluations are presented in Section 4.4.
3.3 Practical Considerations in Sensor Deployment
Data were collected using Zigbee sensors without relying on participants’ Wi-Fi, ensuring the setup was independent of home resources and did not interfere with personal connectivity. Overall, the gateway, Home Assistant and InfluxDB performed reliably. In three houses, a motion sensor malfunction was observed (i.e. one malfunction per house). The sensors went offline within to hours after installation. The data collected were retained, but the three motion sensors were excluded from the analyses. A similar minor challenge arose when one participant repositioned the body and magnet parts of two separate contact sensors too far apart, leaving these sensors permanently open. These experiences highlight the importance of planning for connectivity and on-site contingencies, and underscore the need for a degree of sensor redundancy (e.g. cross-sensor checking) to mitigate potential failures. Additionally, our sensor data have inherent interpretational limitations; for instance, we cannot guarantee that events such as opening cutlery drawers, cabinets, or the refrigerator correspond directly to food preparation or consumption activities.
@@ 4.1.1 Information from participants
The data generated from the interviews allow us to learn about each individual’s ADLs, personal preferences and living context, which, in turn, help us model the properties that reflect general and specific ADLs. Framework Analysis [35], a qualitative analysis method, is applied to reveal patterns and themes in the interview data using NVivo software (Limivero, v14) [36]. Properties of ADLs are identified by considering both our questions during interviews and the routines participants highlighted as important. These properties are then categorised as general (common to everyone, like eating or washing), individual-specific (like pet care or hobbies), or routine but non-daily activities that impact health (like using medication). This approach allows us to have a detailed representation of the properties used to verify general and specific ADLs.
@@ 4.1.2 Sensors
A contact sensor broadcasts its current state every time it changes from open to closed, closed to open or when there is a fluctuation of in the device’s temperature. Repetitions of the same state are taken to be noise in the data triggered by temperature fluctuations so we remove consecutive updates where the state does not change and consider only the first occurrence. This provides a more accurate representation of the participant’s interaction.
A motion sensor broadcasts its state whenever there is a transition between clear and detected states. If the state is detected and no additional motion is registered within seconds, the sensor broadcasts an artificial clear state after the seconds. All clear states are removed from the data as they do not reflect the actual cessation of motion. This provides a more accurate representation of movement, treating motion sensor events as instantaneous. In addition, if continuous clear and detected states occur for a motion sensor, with no changes in other sensors, only the initial detected state of the motion sensor is considered and subsequent states are discarded.
@@ 4.2 Model generator
NuSMV is used to build a model that represents the home and behaviour of the participant. The model is a transition system where changes in sensors can occur one at a time. This is in line with what we observe in the data. The model consists of: i) states, which represent observed configurations of sensor states and where the participant is during the execution of the system; ii) transitions, that correspond to events, which describe how the system evolves from one state to another; and iii) the initial conditions that define the sensors’ starting configuration. The model is verified against a set of logically-encoded properties (see Section 4.3).
@@
| Motion | Hallway | Facing hallway | MotionHallway |
@@
The behaviour of each sensor in the model is encapsulated in a module. An excerpt of the NuSMV code for the contact module is shown in Listing 1. For each contact sensor three variables are defined: its state, activation and deactivation, which are initialised as FALSE (see lines - ). In the module, the state of a contact sensor is a boolean variable where FALSE indicates the body and magnet of the sensor are nearby (e.g. drawers closed as shown in Figure 3b) and TRUE indicates they are apart. The activated boolean variable is set to TRUE at the time step when the state transitions from FALSE ( state) to TRUE (state), as shown in line , and it remains TRUE for one time step only. Note that while activated represents interaction with the person and thus serves as evidence of their activities, the state of a sensor being open – which can last for more than one time step – does not. The deactivated boolean variable is set to TRUE at the time step when the state transitions in the opposite direction, as shown in line , also remaining TRUE for one time step only. Similarly, while deactivated represents interaction by a participant, the state of a sensor being closed does not.
@@
The model is provided with a subtrace of events captured by the different sensors. This subtrace represents the observed behaviour of a participant during a specific period of time. Consider a trace for a set of sensors . Given a pair of unique time markers, and – manually inserted to mark the start and end of a time interval – the subtrace is the set of all events, chronologically ordered, from to , where denotes the time of an event: